# Symbolic Regression with a Learned Concept Library

### v2

@@ 4.6 Using LaSR to discover LLM Scaling Laws
Qualitative Evaluation: Equation 4 describes an empirical relationship between training hyperparameters (training steps) and inference hyperparameters (number of shots). It asserts that increasing the number of shots exponentially increases the model’s performance for low-resource models, while having diminishing gains as the number of training steps of the model increase. This observation is consistent with work in scaling test-time compute [46]. However, if the number of shots is an even number, the exponentially increasing trend reflects over the line $\texttt{score}=E$ and turns into an exponentially decreasing trend. A manual inspection of the training data reveals the majority of samples had an odd number of shots (1, 3, …) which causes this generalization error.
@@
Equation 4 $0.03655\pm 0.00281$ 3
Chinchilla [23] $0.03664\pm 0.00283$ 5
Modified Chinchilla $\mathbf{0.03655\pm 0.00280}$ 5
@@ A.3.3 Hyperparameters
xFigure 7 showcases the hyperparameters used for all our experiments. Wherever possible, we use the default PySR parameters. Additionally, LaSR introduces three new hyperparameters: (1) % of LLM calls, (2) List of user hints, and (3) a dictionary of parameters pertaining to backend LLM communication. Following other methods in SRBench, we utilize only a subset of the necessary operators for solving the Feynman equations, excluding special operators like $\arcsin$ and $\arctan$. These operators are seldom required, and removing them speeds up the search process. We generally set the number of iterations to $40$. However, certain experiments may demand more or less iterations.
@@ A.5.1 Asymmetric comparison with PySR
A common concern with evaluating genetic optimization algorithms w.r.t number of iterations is that the ’hard stop’ after a certain number of iterations might yield populations that haven’t fully converged to their optimal values.
To account for this discrepancy in performance, we conduct an asymmetric comparison with PySR on the Feynman Equations dataset. Specifically, we allow PySR to run uninterrupted for 10 hours per equation, with the maximum number of iterations set to $10^{6}$. We compare the results with that of LaSR run interrupted for 40 iterations per equation (hence the asymmetric comparison). The results are detailed in Table 5.
Overall, we find that, despite running PySR substantially longer than LaSR, LaSR still substantially outperforms PySR. This is because PySR, like other evolutionary algorithms, is susceptible to falling in local minima and converges to this local minima extremely fast. This is why supplementing “local” search strategies with LLM guidance is useful.
@@
A.5.2 Qualitative comparision of Synthetic Dataset equations
@@
GT Equation Predicted Equation $\exp\left(\frac{C-\log(y_{2})}{C-y_{4}}+\left(\sqrt{y_{1}+y_{4}}+\sqrt{y_{2}}\right)\right)$ $\left(\frac{(y_{4}+C+y_{1})}{C}\cdot(y_{2}+C)\right)^{C}+y_{1}$ $\frac{(y_{1}+C)+\left(\sqrt{y_{3}}\cdot\left(y_{1}\cdot\left(C-y_{3}\right)\cdot y_{3}\right)\right)}{\cos(\cos(y_{1}))}$ $y_{3}\cdot\frac{C-\left(y_{1}\cdot\left(y_{3}-\left(\cos\left((y_{1}+C)\cdot C\right)\cdot C\right)+C\right)\right)}{C}-C$ $\sqrt{\exp(y_{1})\cdot y_{1}}\cdot\left(\cos(y_{1})+2y_{1}\right)$ $\left(y_{1}^{C}+C\right)\cdot\left(\sqrt{\exp(y_{1})}-C\right)+\left(C-\left(\frac{\left(y_{1}^{2}\cdot\exp(\cos(y_{1}))\right)}{C}+y_{4}-C\right)\right)$ $\exp\left(\log(y_{4}+y_{2}\cdot y_{1})\right)\cdot\left(\cos(y_{2})-\sqrt{y_{1}}+\frac{y_{3}}{C}-C\right)$ $y_{1}\cdot\left((y_{2}\cdot\cos(y_{2}))-y_{4}-(y_{3}+1)\cdot y_{2}\right)-1$ $(y_{3}+y_{2})\cdot\left(\left((y_{3}+C)\cdot y_{2}+y_{1}+y_{3}\right)-C\right)$ $\left((y_{2}+y_{3})\cdot\left(y_{3}+\left(y_{2}\cdot\left(y_{3}+1\right)\right)+1+y_{1}\right)-C+y_{3}\right)$ $\left(\sqrt{y_{2}}+\exp\left(C+\sqrt{y_{2}}\right)\right)\cdot\left((y_{4}\cdot y_{3})\cdot\log(y_{3})\right)$ $\frac{\left(\left(\sqrt{y_{2}^{C}}-C\right)\cdot(y_{4}\cdot C\cdot(y_{3}-C))-\sin\left(y_{2}^{C}+C\right)\right)\cdot\log(y_{3})}{C}+C$ $\frac{(C\cdot y_{3})\cdot\exp(\sqrt{y_{2}}-y_{1})}{C/(y_{2}+y_{1})}$ $\left(y_{2}\cdot\frac{C-\sqrt{y_{2}^{C}}}{\left(\sqrt{y_{1}^{C}}+C\right)^{y_{1}}}\right)\cdot y_{3}-C$ $\frac{\left(y_{2}+C\right)\cdot\sqrt{y_{1}}}{C}\cdot\sqrt{\frac{\exp(y_{1})}{y_{1}}}$ $\frac{C-y_{2}}{\sin\left(\frac{C}{y_{1}+C}\right)^{C}}-\cos(C-y_{1})$ $y_{1}\cdot\left(\frac{y_{1}^{2}}{\frac{C}{\cos\left(C-y_{4}-C\right)}}\right)/\exp(C\cdot y_{5})$ $\left(\left(y_{1}-C\right)\cdot\left(\sin(y_{4}+C)\cdot y_{1}\right)\cdot(y_{5}+y_{1}-C)\right)\cdot C-C$ $\exp(\log(y_{3})\cdot\sqrt{y_{2}})-\exp\left(\sqrt{(y_{3}+y_{4})+(y_{1}\cdot y_{2})}\cdot C\right)$ $\left(\frac{y_{3}\cdot(y_{2}/C)^{C}+C-y_{4}-y_{2}}{-y_{2}}\right)$
@@
A.5.3 Stochasticity of LaSR and PySR
A common concern with using LLM’s for scientific discovery is that the equations could be obtained due to dataset memorization rather than pure reasoning and learning on the equations form. To explore this further, we run LaSR with the same hyperparameters twice (with different seeds). If the LLM purely relies on memorization, we expect the model to find the same equation form in both experiments. We present results in Table 7.
@@ A.5.3 Stochasticity of LaSR and PySR
Table 7: Performance of LaSR on successive runs with the same hyperparameters on Equation 53 in the Feynman equations dataset (A.5.3). This equation defines the heat flow from a point source (an inverse square law with ground truth formulation: $h=\frac{P}{4\pi R^{2}}$). We evaluate LaSR by running it twice with identical hyperparameters but different random seeds to assess whether the model has memorized the equation’s form. In both runs, LaSR consistently discovers high-performing solutions. Moreover, the resulting functional forms show significant variation between runs, supporting the hypothesis that LaSR’s success is not rooted in memorization.
@@ A.6 Using LaSR to find an LLM Scaling Law
Traditionally, to identify an LLM scaling law, practitioners must first manually posit a “skeleton equation” with a fixed set of known variables and unknown free parameters, and then optimize the unknown parameters based on a dataset of model hyperparameters and resulting dataset fitness [23, 1, 5]. Instead of starting with a predefined equation, we use LaSR to discover the skeleton equation that best fits various subsets of the BigBench dataset. Removing the need to manually posit a skeleton equation simplifies the methodology for finding scaling laws in many ways. First, it removes human preconceptions about the expected relationships between hyperparameters. Second, it increases the number of variables and the type of variables human practitioners can jointly reason about. Finally, it enables positing equations of much higher complexity and variable interdependence than otherwise possible.
@@
We run LaSR with 3840 populations (parallelized across 16 cores) with each population evolving 200 candidates evaluated with the Zygote autodifferentiation backend. We ran LaSR overnight (for 7 hours). BigBench necessitates optimizing a matrix of parameters rather than singular scalar constants. This shifts the compute bottleneck from generating a pool of candidates (in which LaSR is slower than PySR) to evaluating a pool of candidates (which is equally slow for both algorithms).
@@ Results:
score $=\frac{-0.015540}{\left(\frac{\texttt{train\_steps}}{-127140.499}\right)^{\texttt{\#shots}}}+0.363219$ (5)
Where score is the MCQ grade, train_steps is the number of training steps for the model, and #shots is the number of in-context examples provided during inference.
@@ A.8 Further Qualitative Analysis
A.9 Subset of equations discovered by LaSR
[…]
Table 8: Seven equations discovered by LaSR on the Feynman Equations dataset (over PySR). The equations are presented in the form discovered by LaSR, and usually reduce to the ground truth equations after some simplification steps. Note that there are minor discrepancies in the variable names between the ground truth equations in the online lectures (https://www.feynmanlectures.caltech.edu) and the equations in our Feynman equations dataset.

### v3

@@ 4.6 Using LaSR to discover LLM Scaling Laws
Qualitative Evaluation: Equation 4 describes an empirical relationship between training hyperparameters (training steps) and inference hyperparameters (number of shots). It asserts that increasing the number of shots exponentially increases the model’s performance for low-resource models, while having diminishing gains as the number of training steps of the model increase. This observation is consistent with work in scaling test-time compute [46].
@@
Equation 4 $0.03598\pm 0.00265$ 3
Chinchilla [23] $0.03610\pm 0.00268$ 5
Modified Chinchilla $\mathbf{0.03598\pm 0.00264}$ 5
@@ A.3.3 Hyperparameters
Figure 7 showcases the hyperparameters used for all our experiments. Wherever possible, we use the default PySR parameters. Additionally, LaSR introduces three new hyperparameters: (1) % of LLM calls, (2) List of user hints, and (3) a dictionary of parameters pertaining to backend LLM communication. Following other methods in SRBench, we utilize only a subset of the necessary operators for solving the Feynman equations, excluding special operators like $\arcsin$ and $\arctan$. These operators are seldom required, and removing them speeds up the search process. We generally set the number of iterations to $40$. However, certain experiments may demand more or less iterations.
@@ A.5.1 Asymmetric comparison with PySR
A common concern with evaluating genetic optimization algorithms w.r.t number of iterations is that the ‘hard stop’ after a certain number of iterations might yield populations that haven’t fully converged to their optimal values.
To account for this discrepancy in performance, we conduct an asymmetric comparison with PySR on the Feynman Equations dataset. Specifically, we allow PySR to run uninterrupted for 10 hours per equation, with the maximum number of iterations set to $1\times 10^{6}$. We compare the results with that of LaSR run interrupted for 40 iterations per equation (hence the asymmetric comparison). The results are detailed in Table 5.
Overall, we find that, despite running PySR substantially longer than LaSR, LaSR still substantially outperforms PySR. This is because PySR, like other evolutionary algorithms, is susceptible to falling in local minima and converges to this local minima extremely fast. This indicates that supplementing ‘local search’ strategies with LLM guidance is useful for symbolic regression.
@@
A.5.2 Subset of equations discovered by LaSR on the Feynman Dataset
[…]
Table 6: A subset of equations discovered by LaSR on the Feynman Equations dataset (over PySR). The equations are presented in the form discovered by LaSR, and usually reduce to the ground truth equations after some simplification steps. Note that there are minor discrepancies in the variable names between the ground truth equations found in the online Feynman Lectures (https://www.feynmanlectures.caltech.edu) and those in our Feynman Equations dataset.
A.5.3 Qualitative comparison of Synthetic Dataset equations
@@
Ground Truth Equation Discovered Equation $\exp\left(\frac{C-\log(y_{2})}{C-y_{4}}+\left(\sqrt{y_{1}+y_{4}}+\sqrt{y_{2}}\right)\right)$ $\left(\frac{(y_{4}+C+y_{1})}{C}\cdot(y_{2}+C)\right)^{C}+y_{1}$ $\frac{(y_{1}+C)+\left(\sqrt{y_{3}}\cdot\left(y_{1}\cdot\left(C-y_{3}\right)\cdot y_{3}\right)\right)}{\cos(\cos(y_{1}))}$ $y_{3}\cdot\frac{C-\left(y_{1}\cdot\left(y_{3}-\left(\cos\left((y_{1}+C)\cdot C\right)\cdot C\right)+C\right)\right)}{C}-C$ $\sqrt{\exp(y_{1})\cdot y_{1}}\cdot\left(\cos(y_{1})+2y_{1}\right)$ $\left(y_{1}^{C}+C\right)\cdot\left(\sqrt{\exp(y_{1})}-C\right)+\left(C-\left(\frac{\left(y_{1}^{2}\cdot\exp(\cos(y_{1}))\right)}{C}+y_{4}-C\right)\right)$ $\exp\left(\log(y_{4}+y_{2}\cdot y_{1})\right)\cdot\left(\cos(y_{2})-\sqrt{y_{1}}+\frac{y_{3}}{C}-C\right)$ $y_{1}\cdot\left((y_{2}\cdot\cos(y_{2}))-y_{4}-(y_{3}+1)\cdot y_{2}\right)-1$ $(y_{3}+y_{2})\cdot\left(\left((y_{3}+C)\cdot y_{2}+y_{1}+y_{3}\right)-C\right)$ $\left((y_{2}+y_{3})\cdot\left(y_{3}+\left(y_{2}\cdot\left(y_{3}+1\right)\right)+1+y_{1}\right)-C+y_{3}\right)$ $\left(\sqrt{y_{2}}+\exp\left(C+\sqrt{y_{2}}\right)\right)\cdot\left((y_{4}\cdot y_{3})\cdot\log(y_{3})\right)$ $\frac{\left(\left(\sqrt{y_{2}^{C}}-C\right)\cdot(y_{4}\cdot C\cdot(y_{3}-C))-\sin\left(y_{2}^{C}+C\right)\right)\cdot\log(y_{3})}{C}+C$ $\frac{(C\cdot y_{3})\cdot\exp(\sqrt{y_{2}}-y_{1})}{C/(y_{2}+y_{1})}$ $\left(y_{2}\cdot\frac{C-\sqrt{y_{2}^{C}}}{\left(\sqrt{y_{1}^{C}}+C\right)^{y_{1}}}\right)\cdot y_{3}-C$ $\frac{\left(y_{2}+C\right)\cdot\sqrt{y_{1}}}{C}\cdot\sqrt{\frac{\exp(y_{1})}{y_{1}}}$ $\frac{C-y_{2}}{\sin\left(\frac{C}{y_{1}+C}\right)^{C}}-\cos(C-y_{1})$ $y_{1}\cdot\left(\frac{y_{1}^{2}}{\frac{C}{\cos\left(C-y_{4}-C\right)}}\right)/\exp(C\cdot y_{5})$ $\left(\left(y_{1}-C\right)\cdot\left(\sin(y_{4}+C)\cdot y_{1}\right)\cdot(y_{5}+y_{1}-C)\right)\cdot C-C$ $\exp(\log(y_{3})\cdot\sqrt{y_{2}})-\exp\left(\sqrt{(y_{3}+y_{4})+(y_{1}\cdot y_{2})}\cdot C\right)$ $\left(\frac{y_{3}\cdot(y_{2}/C)^{C}+C-y_{4}-y_{2}}{-y_{2}}\right)$
@@
A.5.4 Stochasticity of LaSR and PySR
A common concern with using LLM’s for scientific discovery is that the equations could be obtained due to dataset memorization rather than reasoning and learning on the equations form. To explore this further, we run LaSR with the same hyperparameters twice with different seeds. We expect LaSR to behave like PySR (or any stochastic evolutionary algorithm) and produce two syntactically different programs that achieve a similar data fitness score. However, if the algorithm purely relies on training set memorization, we expect the model to find the same equation form in both experiments. We present results in Table 8.
@@ A.5.4 Stochasticity of LaSR and PySR
Table 8: Performance of LaSR on successive runs with the same hyperparameters on Equation 53 in the Feynman equations dataset (A.5.4). This equation defines the heat flow from a point source (an inverse square law with ground truth formulation: $h=\frac{P}{4\pi R^{2}}$). We evaluate LaSR by running it twice with identical hyperparameters but different random seeds to assess whether the model has memorized the equation’s form. In both runs, LaSR consistently discovers high-performing solutions. Moreover, the resulting functional forms show significant variation between runs, supporting the hypothesis that LaSR’s performance is not rooted in memorization.
@@ A.6 Using LaSR to find an LLM Scaling Law
Traditionally, to identify an LLM scaling law, practitioners must first manually posit a skeleton equation with a fixed set of known variables and unknown free parameters, and then optimize the unknown parameters based on a dataset of model hyperparameters and resulting dataset fitness [23, 1, 5]. Instead of starting with a predefined equation, we use LaSR to discover the skeleton equation that best fits various subsets of the BigBench dataset. Removing the need to manually posit a skeleton equation simplifies the methodology for finding scaling laws in many ways. First, it removes human preconceptions about the expected relationships between hyperparameters. Second, it increases the number of variables and the type of variables human practitioners can jointly reason about. Finally, it enables positing equations of much higher complexity and variable interdependence than otherwise possible.
@@
We run LaSR with 3840 populations (parallelized across 16 cores) with each population evolving 200 candidates evaluated with the Zygote autodifferentiation backend. We ran LaSR overnight (for 7 hours). BigBench necessitates optimizing a matrix of parameters rather than singular scalar constants. This shifts the compute bottleneck from generating a pool of candidates (in which LaSR is slower than PySR) to evaluating a pool of candidates (which is equally slow for both algorithms). We also impose regularization constraints to accelerate the optimization procedure and avoid degenerate equations. Specifically, we expect a monotonically increasing performance trend with respect to the number of training steps.
@@ Results:
score $=\frac{-0.0248235}{\left(\frac{\texttt{train\_steps}}{116050.999}\right)^{\texttt{\#shots}}}+0.360124$ (5)
Where score is the MCQ grade, train_steps is the number of training steps for the model, and #shots is the number of in-context examples provided during inference. LaSR maintains a Pareto frontier of the simplicity-accuracy tradeoff between various best performing programs. We present all the programs in the pareto frontier of this experiment in Table 9.
Complexity Equation
2 $A$
4 $A-B$
6 $A+\left(\frac{-C}{A}\right)$
8 $A+\left(\frac{-D}{\texttt{training\_steps}^{\texttt{number\_of\_shots}}}\right)$
10 $A+\left(\frac{-D}{\left(\frac{\texttt{training\_steps}}{B}\right)^{\texttt{number\_of\_shots}}}\right)$
13 $\frac{A+E}{\texttt{training\_steps}+\frac{\cos(\texttt{training\_steps})}{\cos(\texttt{training\_steps})+F}}$
16 $\frac{A+G}{\left(\frac{H}{I+\cos(\texttt{training\_steps})}\cdot J\right)+(\texttt{training\_steps}-K)}$
17 $\frac{E+A}{\texttt{training\_steps}+\frac{\cos(\texttt{training\_steps})}{\cos(\texttt{training\_steps})+(F-L\cdot B)}}$
19 $\frac{A+M}{\left(\frac{N+A}{O+\cos(\texttt{training\_steps})}\cdot P\right)+(\texttt{training\_steps}-Q)}$
21 $\frac{A+M}{\left(\frac{N+\cos(A+\texttt{total\_params})}{O+\cos(\texttt{training\_steps})}\cdot P\right)+(\texttt{training\_steps}-Q)}$
25 $\frac{A+G}{\left(\texttt{training\_steps}-K+\left(\frac{H+\cos(A+\texttt{non\_embedding\_params})}{R+\cos(\texttt{training\_steps})}\cdot J-S\right)\right)\cdot T}$
26 $\frac{A+G}{\left(\texttt{training\_steps}-K+\left(\frac{H+\cos(A+\texttt{non\_embedding\_params})}{R+\cos(\texttt{training\_steps})}\cdot J-S\right)\right)\cdot T}$
28 $\frac{A+G}{\left(\texttt{training\_steps}-K+\left(\frac{H+B+\cos(A+\texttt{total\_params})}{I+\cos(\texttt{training\_steps})}\cdot J-S\right)\right)\cdot T}$
30 $\frac{A+G}{\left(\texttt{training\_steps}-K+\left(\frac{H+B+\cos(A+\texttt{total\_params}^{U})}{I+\cos(\texttt{training\_steps})}\cdot J-S\right)\right)\cdot T}$
32 $\frac{(A+G)-V}{\left(\texttt{training\_steps}-K+\left(\frac{\cos(A+\texttt{total\_params}^{W})+H+B}{I+\cos(\texttt{training\_steps})}\cdot J-S\right)\right)\cdot T}$
Table 9: Pareto frontier of LaSR for the BigBench experiment (Section A.6). Equation 4 is shown as the equation with complexity $10$ in this table.